Mole fraction of a solute in benzene(C6H6) is 0.2 then find molarity of solute

Grade:12th Pass

1 Answers

Arun
5 years ago
Dear student
 
mole fraction = 0.2.therefore in 100 moles, 20 moles are of solute and 80 moles are of benzene(solvent)mass of benzene = 80 * Mol.Wt of Benzene = 80* 78 = 6240gfor 6240 g of solvent we have 20 moles of soluteTherefore for 1000g of solvent we have (20)*(1000)/(6240) = 3.20 molal
 
Regards
Arun askIITians forum expert()